OpenStudy (anonymous):

show that equation: a = V^2_(av) / R is dimensionally correct. i.e. show that the units match on both sides of the equation

OpenStudy (anonymous):

[a]=L.T^-2 [v]=L.T^-1 [R]=L Hence [v^2/R]=[v]^2/[R]=L^2.T^-2/L=L.T^-2=[a]

OpenStudy (anonymous):

where L=''unit of length'', T=''unit of time''

OpenStudy (anonymous):

and [x]=''dimensions of x''

OpenStudy (anonymous):

thank you!